### v1

# Multimodal Foundation Models for Early Disease Detection

Abstract
Healthcare generates diverse streams of data, including electronic health records (EHR), medical imaging, genetics, and ongoing monitoring from wearable devices. Traditional diagnostic models frequently analyze these sources in isolation, which constrains their capacity to identify cross-modal correlations essential for early disease diagnosis. Our research presents a multimodal foundation model that consolidates diverse patient data through an attention-based transformer framework. At first, dedicated encoders put each modality into a shared latent space. Then, they combine them using multi-head attention and residual normalization. The architecture is made for pretraining on many tasks, which makes it easy to adapt to new diseases and datasets with little extra work. We provide an experimental strategy that uses benchmark datasets in oncology, cardiology, and neurology, with the goal of testing early detection tasks. The framework includes data governance and model management tools in addition to technological performance to improve transparency, reliability, and clinical interpretability. The suggested method works toward a single foundation model for precision diagnostics, which could improve the accuracy of predictions and help doctors make decisions.
I Introduction
The swift aggregation of manifold patient data poses both considerable analytical hurdles and unique opportunities. Healthcare data today come in many modalities, ranging from imaging such as MRI and CT scan, to sequential streams like wearable sensor outputs and electronic health records (EHRs). They also include audio (heart sounds, breathing patterns, or interview recordings), written sources (clinical documentation and scientific articles), video (for example, surgical recordings), as well as molecular-level information such as genomics and proteomics [1]. EHR further capture a timeline of clinical events, where variables such as diagnoses, medications, and procedures are captured at each visit . When combined, these sequences provide a detailed record of an individual patient’s medical history across multiple encounters [2] [3].
As healthcare providers and researchers aim to improve patient care through precision medicine [4], clinical decision-making significantly depends on the synthesis of these information sources. However, current predictive models used for this purpose are typically limited to a single data modality, which constrains their ability to provide comprehensive, patient-centric predictions as they could not catch the intricate interdependencies that exist across these heterogeneous data types. This limitation reduces the potential for genuinely patient-centric insights and constrains predictive accuracy.
In recent years, as deep learning has been producing models and tools capable of capturing complex relationships across diverse data modalities, attention-based architectures in particular, have shown the ability to dynamically evaluate the relevance of different inputs, allowing models to emphasize the information most critical for a given patient or predictive task. When pretrained on large, heterogeneous datasets, such models learn generalizable representations that can be adapted to multiple downstream diagnostic applications [5]. This paradigm, well established in natural language processing and computer vision [6], positions our approach as a foundation model for healthcare, providing reusable multimodal representations to support early disease detection and other clinical tasks. Despite these advancements, there are still many challenges because multimodal datasets are often fragmented, and individual patients may have partial or missing modalities [7]. Additionally, high-dimensional data, like whole-genome sequences or high-resolution images, make model integration and training even more difficult. Furthermore, models must provide transparent reasoning to support physician trust and facilitate actionable decisions; interpretability is a crucial requirement for clinical adoption.
The attention-based transformer architecture has proven highly effective for large-scale pretraining, where it learns rich contextual representations that enhance performance on a variety of downstream tasks, including sentiment analysis, information retrieval, and entity recognition [8]. Building on this paradigm, foundation models trained on massive and diverse datasets extend these benefits by enabling stronger contextual reasoning, broader generalization, and emergent prompting capabilities during inference [9]. That is why we propose a transformer-based foundation model architecture to resolve user challenges, which is intended to integrate EHR, imaging, genomic, and wearable sensor data. Our method utilizes cross-modal attention mechanisms to identify interdependencies between modalities and to facilitate pretraining on a variety of tasks, thereby facilitating the development of generalizable, patient-specific representations. This positions our approach as a multimodal foundation model, capable of serving as a base for diverse clinical prediction and diagnostic tasks. This work endeavors to facilitate the advancement of integrated diagnostic tools that are capable of enhancing patient outcomes and supporting precision medicine initiatives by integrating heterogeneous clinical data streams within a unified computational framework.
II Related Work
II-A Multimodal Integration
The wide use of EHR systems has led to the development of predictive models that can support clinical care [2]. As healthcare data now extend far beyond EHRs to include imaging, genomics, and signals from wearable devices, these models are increasingly applied to improve patient outcomes [10]. However, most current methods remain limited to a single data type; so they miss important cross-modal information, struggle with incomplete inputs, and often rely on specialized fusion methods that do not scale well to high-dimensional biomedical data. In practice, clinical information is multimodal. Integrating different sources of data is often necessary for accurate diagnosis and effective treatment [1]. Studies show that combining EHRs with imaging or genomics can improve diagnostic accuracy when paired with deep learning methods [11]. At the same time, foundation models—large pretrained systems designed to adapt across many tasks [6] —are becoming increasingly important in healthcare for their ability to generalize across data types and domains [9],[12]. Together, these developments suggest a natural convergence: Multimodal foundation models provide a way to unify heterogeneous clinical data within a single framework, with the potential to improve early disease detection and enable more precise, patient-centered care.
II-B Attention-Based Models
Attention-based transformer models came out in 2017 [13] and quickly gained popularity for working with sequential data. They handle sequences differently from older models like Recurrent Neural Networks (RNNs), which process data step by step. Instead, transformers can consider the whole sequence at once, which tends to make it easier to spot patterns that span long sections of the data [14]. In addition, they scale fairly well when the dataset is large and can be trained in parallel. That combination of flexibility and efficiency is one reason researchers have started using them in many areas, including healthcare [15].
In healthcare, transformers have been applied to handle large and messy data sources, especially clinical notes and records. They can highlight which pieces of information matter most for a given prediction, making them useful for tasks where not every variable has the same importance [16]. This property is also what makes them appealing for multimodal learning: the model can weigh structured EHR entries against imaging, genomic profiles, or even wearable sensor data without relying on heavy feature engineering [17].
Even so, current multimodal healthcare systems are usually limited. For instance, most models concentrate on single inputs, and leave out other valuable sources like wearables or genomics. Another limitation is that few of them use foundation-style pretraining, where the model is trained broadly across tasks and then reused for specific problems. Without this, adaptability is reduced and each new task often needs a separate model.
This is the gap we target in our work. By combining multimodal integration with foundation-model pretraining, we aim to create patient representations that are broad enough to transfer between tasks yet sensitive to early signals of disease that might otherwise go unnoticed.
III Proposed Framework
We propose a multimodal transformer-based framework to combine diverse biomedical data sources for early illness diagnosis. Unlike unimodal techniques, which handle only one type of input (e.g., EHR or imaging), our architecture combines a variety of clinical data streams, such as electronic health records, medical imaging, genomic sequences, and wearable sensor data. The architecture stresses both flexibility and robustness, allowing predictions even when some modalities are absent, as well as extensibility to new patient data sources that were not available during training. Figure 1 shows a patient-centric multimodal transformer design.
Patient DataModality EncodersEHR, Imaging, Genomics, WearablesFeature EmbeddingModality-specific EncodingsCross-Modal Attention FusionMulti-Head Attention + Modality DropoutNormalization & ResidualsLayer Norm, Residual ConnectionsDisease PredictionEarly Detection, Risk Scores Data Governance Audit Trails Missing Data Handling Normalization Standardization Model Management Training Logs Performance Metrics Interpretability
Figure 1: Multimodal transformer architecture
III-A Problem Definition
Let each patient record be denoted by:
$\mathcal{P}_{i}=\{X^{ehr}_{i},X^{img}_{i},X^{gen}_{i},X^{sens}_{i}\},$
where $X^{ehr}_{i}\in\mathbb{R}^{T\times d_{ehr}}$ represents temporal EHR data, $X^{img}_{i}\in\mathbb{R}^{H\times W\times C}$ corresponds to imaging modalities, $X^{gen}_{i}\in\mathbb{R}^{L_{gen}\times d_{gen}}$ denotes genomic features, and $X^{sens}_{i}\in\mathbb{R}^{T_{s}\times d_{sens}}$ encodes wearable sensor signals.
The predictive objective is:
$\hat{y}_{i}=f_{\theta}(\mathcal{P}_{i}),\quad y_{i}\in\{0,1\}^{K},$
where $\hat{y}_{i}$ denotes the probability distribution over $K$ disease classes.
As not all patients will have complete data across all modalities, our framework is designed to accommodate incomplete records by incorporating modality dropout during training as well as enabling inference on any available subset of $\{X^{ehr},X^{img},X^{gen},X^{sens}\}$, which ensures practical applicability in clinical settings where data coverage is uneven.
III-B Modality-Specific Encoders
Each of the modality is transformed into a latent representation by a dedicated encoder:
$h^{m}_{i}=\phi_{m}(X^{m}_{i}),\quad m\in\{ehr,img,gen,sens\}.$
Input: Patient data $\mathcal{P}=\{X^{ehr},X^{img},X^{gen},X^{sens}\}$
Output: Latent representations $\{h^{ehr},h^{img},h^{gen},h^{sens}\}$
foreach modality $m\in\{ehr,img,gen,sens\}$ do
 $h^{m}\leftarrow\phi_{m}(X^{m})$ ;
 // Apply modality-specific encoder
return $\{h^{ehr},h^{img},h^{gen},h^{sens}\}$
Algorithm 1 Modality Encoding
Here, raw inputs are structured into feature spaces. For EHR, sequential embeddings capture temporal dependencie, for imaging data are processed through Convolutional Neural Networks (CNNs) or Vision Transformers (ViTs) to extract spatial hierarchies, genomics are encoded with sequence models such as 1D CNNs, while wearable signals are modeled with temporal CNNs or Gated Recurrent Units (GRUs). By design, each of the encoder learns modality-specific representations $h^{m}$ that preserve essential features while normalizing heterogeneous input types.
III-C Cross-Modal Attention Fusion
Encoded features are aggregated to form a unified embedding. Let:
$Z=\{h^{ehr}_{i},h^{img}_{i},h^{gen}_{i},h^{sens}_{i}\}.$
Cross-modal attention operates over this set:
$\text{Attention}(Q,K,V)=\text{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V,$
with
$Q=W_{Q}Z,\quad K=W_{K}Z,\quad V=W_{V}Z.$
The fused embedding is:
$h^{fusion}_{i}=\text{Concat}(\text{head}_{1},\dots,\text{head}_{H})W_{O}.$
Input: Encoded representations $Z=\{h^{ehr},h^{img},h^{gen},h^{sens}\}$
Output: Fused embedding $h^{fusion}$
$Q\leftarrow W_{Q}Z,\;K\leftarrow W_{K}Z,\;V\leftarrow W_{V}Z$;
for head =1=1 to $H$ do
 $head_{h}\leftarrow\text{softmax}\!\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$;
$h^{fusion}\leftarrow\text{Concat}(head_{1},\dots,head_{H})W_{O}$;
return $h^{fusion}$
Algorithm 2 Cross-Modal Fusion
Here representations across modalities are integrated and queries, keys, and values are projected from the modality embeddings. Multi-head attention ensures that the model can capture distinct relationships (e.g., correlating imaging abnormalities with lab results or linking genetic variants to wearable data patterns). The concatenation of attention heads followed by a linear projection yields $h^{fusion}$, a joint patient embedding suitable for downstream classification.
III-D Training Strategy
The training protocol consists of two stages: first, a large corpus of multimodal data is used for self-supervised pretraining and second, the pretrained model is then fine-tuned with supervised labels for disease prediction. Objectives of the training phase include: (i) masked reconstruction of missing inputs and (ii) cross-modal contrastive learning to align paired modalities.
Input: Unlabeled data $\mathcal{D}_{pre}$, labeled data $\mathcal{D}_{task}$
Output: Optimized parameters $\theta^{*}$
foreach batch $\mathcal{B}\subset\mathcal{D}_{pre}$ do
   Compute reconstruction loss $\mathcal{L}_{mask}$;
   Compute contrastive loss $\mathcal{L}_{contrast}$;
   Update $\theta$ using $\nabla(\mathcal{L}_{mask}+\alpha\mathcal{L}_{contrast})$;
foreach batch $\mathcal{B}\subset\mathcal{D}_{task}$ do
   Encode and fuse modalities;
   Predict disease label $\hat{y}$;
[…]
   Update $\theta$ with gradient descent;
return $\theta^{*}$
Algorithm 3 Training Procedure
During pretraining, the encoders and fusion module learn general multimodal patterns, which helps the model handle noisy data. On the other hand, masked reconstruction encourages it to fill incomplete information, while contrastive learning aligns embeddings from different modalities. In the fine-tuning stage, the network is tailored to the disease detection task by optimizing cross-entropy loss on labeled data. This two-step process strikes a balance between broad general knowledge and task-specific performance.
III-E End-to-End Reasoning Flow
Input: Patient record $\mathcal{P}=\{X^{ehr},X^{img},X^{gen},X^{sens}\}$ (some modalities may be missing)
Output: Prediction $\hat{y}$, uncertainty $u$, explanation $E$
// Input acquisition & preprocessing
$X^{m}\leftarrow\text{Preprocess}(X^{m})\;\forall m\in\{ehr,img,gen,sens\}$;
Simulate/record missing-modality mask $M$;
// Unimodal representation learning
foreach modality $m$ available do
 $z^{m}\leftarrow$ Encode($X^{m}$) ;
 // encoder: Transformer / CNN / 1D-CNN / TCN
 $e^{m}\leftarrow$ Embed($z^{m}$) ;
 // project to shared latent space
// Cross-modal alignment (optional contrastive step during pretraining)
Align $\{e^{m}\}$ with contrastive or projection losses (pretraining);
// Multimodal reasoning via fusion
$h^{(0)}\leftarrow\text{Aggregate}(\{e^{m}\})$ ;
// e.g., concatenation or learned pooling
for $\ell=1$ to $L$ do
 $h^{(\ell)}\leftarrow$ CrossModalFuse $(h^{(\ell-1)},\{e^{m}\},M)$ ;
 // multi-head cross-modal attention + residuals
   Optional: apply LayerNorm and Feed-Forward block
// Prediction & uncertainty
$\hat{y}\leftarrow$ PredictHead $(h^{(L)})$;
[…]
// e.g., MC-dropout, ensemble, or Bayesian head
// Explanation / Chain-of-Thought trace
$E\leftarrow\text{ExtractAttentionMaps}(\{\text{attention weights from }\ell\})$;
// Feedback / continual update (deployed system)
if feedback available (label/outcome) then
 ApplyFeedback $(\theta,\text{feedback})$ ;
 // fine-tune or federated update
return $\hat{y},u,E$
Algorithm 4 Chain-of-Thought Template: Multimodal Diagnostic Inference
This the chain-of-thought template that depicts the proposed framework’s end-to-end reasoning process to supplement the modality-specific algorithms. This template shows how the whole thing works as a diagnostic workflow. The process starts with raw data preprocessing and modality-specific encoding, then moving on to cross-modal fusion and prediction, uncertainty estimates, explanation creation, and updates driven by feedback.
IV Experimental Design
In order to evaluate the proposed multimodal transformer framework rigorously, we present a comprehensive evaluation strategy that demonstrates clinical relevance, robustness, and feasibility. Although the experiments are conceptual, the design guarantees that the framework can be systematically validated upon its implementation.
IV-A Datasets
The framework is intended to manage a wide range of multimodal patient records. We suggest that an evaluation be conducted on publicly available datasets that encompass complementary types of data across various disease domains:
  • MIMIC-IV: Enables the evaluation of longitudinal modeling and temporal feature extraction by providing structured EHR data, laboratory tests, and clinical notes.
  • UK Biobank: Supports cross-modal integration and generalizability testing across population-level cohorts by incorporating genomic, health record, and imaging data.
  • ADNI: Provides a benchmark for the prediction of neurodegenerative diseases by integrating cognitive assessments, biomarker data, and MRI and PET imaging.
IV-B Baselines and Comparative Approaches
In order to contextualize performance, the proposed framework would be evaluated in comparison to:
  • Unimodal baselines: LSTM/transformer models for EHR sequences, CNN or ViT models for imaging, and gradient-boosted trees for genomics.
  • Multimodal baselines: VAE-based integration, GNN-based patient similarity approaches, and concatenation-based fusion.
IV-C Evaluation Metrics
Evaluation metrics are chosen to indicate the model’s reliability and its clinical significance.
  • AUROC: Evaluates the model’s ability to correctly classify patients based on their specific diseases.
  • AUPRC: Highlights the importance of performance on positive cases to mitigate class imbalance, which is essential for rare conditions.
  • F1-Score and Accuracy: Articulate the quality of classification in a manner that is both clear and understandable.
  • Calibration Error: Assesses the reliability of predicted probabilities, which is essential for clinical decision support.
IV-D Proposed Analyses
To comprehensively validate the framework, we recommend performing a range of analyses:
  • Multimodal vs Unimodal Approaches: To quantify the benefits of integrating multiple data types, each modality will be evaluated separately and in combination.
  • Lack of modality robustness: The model’s capacity to handle incomplete records will be demonstrated by systematically excluding one or more modalities.
  • Ablation Research: The removal of components such as modality-specific encoders, cross-modal attention, or pretraining objectives will explicate their individual contributions.
  • Tests of Generalizability: The potential of transfer learning and its applicability across diseases will be evaluated by utilizing pretrained representations on diverse datasets or patient populations.
V Discussion
Our proposed multimodal transformer framework offers a structured approach to the integration of heterogeneous clinical data in order to facilitate the early detection of diseases. The framework can capture intricate relationships that unimodal models fail to capture by encoding EHRs, imaging, genomics, and wearable signals and fusing them through cross-modal attention. This design facilitates interpretability through attention analysis, robustness to missing data, generalizable patient representations obtained through pretraining, and enhanced predictive performance. The following challenges persist: high computational demands, limited publicly available multimodal datasets, variability and quality of data, and integration into clinical workflows. Future directions include the integration of the framework with real-time clinical decision support systems, the exploration of lightweight transformer variants, the development of continual learning and domain adaptation strategies, and the addition of new data sources. In general, this method provides a scalable, interpretable path to the identification of early diseases that are more precise and personalized.
VI Use Cases in Early Disease Detection
VI-A Oncology: Multimodal Detection of Cancer
The detection of early cancer remains one of the most significant challenges in clinical medicine, as the stage at which the disease is diagnosed has a strong correlation with the outcome. Traditional methods heavily depend on imaging or histopathology; however, these modalities may not be able to detect subtle precancerous or early neoplastic changes when used in isolation. A multimodal foundation model can integrate radiological images, pathology slides, genomic alterations, and structured EHR data, including family history and laboratory results. For instance, longitudinal EHR data may offer patterns of symptom evolution, while genomic biomarkers associated with tumor predisposition could be contextualized with radiographic features of suspicious lesions. The fusion of these complementary signals has the potential to identify malignancies at a preclinical stage, thereby reducing false negatives and enhancing screening specificity. In practice, this system could serve as an auxiliary diagnostic aid, indicating high-risk patients for additional evaluation prior to the complete emergence of clinical symptoms.
VI-B Cardiovascular Disease: Predicting Heart Failure
Despite the fact that cardiovascular disease continues to be one of the most prevalent causes of illness and mortality on a global scale, it remains a significant obstacle to early identification of individuals who are at risk. The majority of contemporary models are predicated on static data from electronic health records, including echocardiographic findings or laboratory results. However, other data including wearable sensor data (heart rate variability, activity levels), imaging from echocardiography or cardiac MRI, and genetic risk scores from sequencing could be combined to further develop a multimodal approach. For instance, imaging indicators of diminished ventricular function and genetic predispositions to arrhythmias could be examined in conjunction with subtle variations in cardiac rhythm detected by wearables. These streams of information could be combined to facilitate real-time monitoring and ongoing risk assessment which will allow clinicians to take preventive measures earlier and reduce hospitalizations and deleterious cardiac events as a result.
VI-C Neurology: Early Detection of Neurodegenerative Disorders
Neurodegenerative diseases like Alzheimer’s and Parkinson’s show up slowly, with early signs of disease appearing years before the symptoms do. The currently used models that only employ scanning or cognitive tests often miss the first signs that a disease is getting worse. But in a multimodal approach, neuroimaging (structural and functional), genomic variants linked to neurodegeneration, longitudinal EHR data recording subtle behavioral or cognitive complaints, and continuous monitoring from wearable devices that track motor activity could all be a apart of. Combining these different signs, the model will able to find changes that were happening behind the scenes before they became noticeable as a loss of cognitive function. For example, sleep problems caused by wearables may work well with early imaging signs of hippocampal atrophy to make a full risk score. This makes it possible for patients to be quickly enrolled in early-stage interventions or clinical studies, which increases the chances of finding treatments that change the course of the disease.
VII Conclusion
We outline a framework for multimodal foundation models aimed at healthcare diagnostics. The idea is simple: instead of relying on one type of data, such as EHRs, the model brings together records, images, genetic information, and signals from wearable devices. A transformer-based setup is used to handle this mix. Each type of input is first processed by its own encoder, so important details are not lost. These are then linked through attention mechanisms, and pretraining helps the model build general structure before it is tuned for specific clinical tasks. The design also stresses aspects that are often overlooked, like data governance, model monitoring, and interpretability. In practice, these matter as much as raw accuracy, since clinicians need to understand decisions, trust outputs, and know that data quality is maintained. The framework structure is meant to be flexible across diseases and patient groups. Over time, we hope this kind of system could allow earlier diagnosis, better risk prediction, and treatment tailored to individuals—steps toward precision medicine.